# Exploring the operational limits of Mirnov coil arrays in stellarators by means of a synthetic diagnostic

### v1

@@ IV.1.1 Poloidal array
Figure 11: Ratio of squared perpendicular $|\delta B_{\circledast}|$ and parallel $|\delta B_{\circledcirc}|$ amplitudes detected by the poloidal coils for $n=0$ modes with different poloidal mode numbers.
@@
Figure 12: 2D projection $\delta B_{\circledcirc}(t)$ of the general $\delta{\mathbf{B}}(t)$ polarization ellipses for an $m=2$, $n=0$ mode. Vectors $\mathbf{P}$ (blue arrows) and the vectors indicating the major (black arrows) and minor (red arrows) axis of the polarization ellipses at each location are shown.
@@
As a first approach, and considering the geometry of the poloidal array, we restrict ourselves to the projection of the general 3D polarization ellipse on the RZ plane, trying to find out if there is an optimal orientation for the coils, always keeping $\mathbf{P}$ in the RZ plane. In figure 11 we represent the ratio between the amplitude of the oscillating field perpendicular ($|\delta B_{\circledast}|$) and parallel ($|\delta B_{\circledcirc}|$) to the RZ plane. A very low value of this ratio indicates that the oscillating field lies mostly in the RZ plane, so we can expect a better performance of the diagnostic in such cases. This ratio depends on both the mode numbers and the radial position of the modes, and there are instances where it is not small. Figure 12 shows the (normalized) projection of the $\delta B_{\circledcirc}(t)$ polarization ellipse with their major and minor axes and the vectors $\mathbf{P}$, that determine the coil orientations, for an $m=2$, $n=0$ mode. Note the changes in angle between $\mathbf{P}$ and the major axis of the ellipse, that become more pronounced on the rightmost coils.
@@
It might seem that knowing the orientation of the polarization ellipse at each point and applying the periodogram technique to the components of $\delta B_{\circledcirc}(t)$ in the local polarization basis defined by the major and minor axis of the measured ellipses could lead to an improvement in the mode identification. However, this method has been tested and it has proved to not be effective, and projecting over the design orientation given by $\mathbf{P}$, which in this case is approximately tangent to the plasma surface, yields much better results. Furthermore, as the polarization is mode-dependent, one cannot design an uni-axial coil array that performs well for any mode, and the experimental use of the aforementioned procedure is not feasible without having tri-axial sensors. Going back to figure 12, we see that the coils where the angle between $\mathbf{P}$ and the major axis of the ellipse changes more rapidly are the ones that are closest to the plasma. It could be advantageous, then, to restrict ourselves to the coils furthest from the plasma, hoping for a decrease in the distortion of the mode structure and an improvement of the detection accuracy. Figure 13 shows the results of the mode analysis with the rightmost five coils removed on each side, and we observe that the detection accuracy does not improve, and off-by-one errors appear more frequently. The detection limits (the radial position and mode width where the mode identification is not possible), on the other hand, do not change significantly. This could be caused by a lack of angular coverage of the plasma column, but that would mainly affect the detection of low-$m$ modes ($m\leq 3$), which does not seem to be the case. Alternatively, the issue may lie in the coils furthest from the plasma, as the part of the plasma that contributes to the Biot-Savart integral with significant intensity increases with coil distance (as the contribution of a volume element is proportional to $1/r^{2}$).
It is natural then to inquire about the effects of coil distance over the measured signal. For that, a set of signals with different mode numbers have been simulated for a hypothetical linear array of coils located at $\varphi=0$, $z=0$ and distributed with increasing radial distance $R$. The amplitude of the perturbed field at the $i$-th coil position is:
$A_{i}\equiv\sqrt{\Bigl{\langle}\absolutevalue{\delta\dot{\mathbf{B}_{i}}}^{2}\Bigr{\rangle}_{t}}$ (10)
where $\langle\cdot\rangle_{t}$ denotes a time average. In the multipolar approximation Jackson (1999), this signal follows an inverse-power law, decaying as $1/r^{m+1}$, so an attempt has been made to fit the field amplitudes to the function:
$A(r)=\dfrac{A_{0}(m)}{(r_{\text{LCFS}}+\delta r)^{m+1}}$ (11)
Where $A_{0}(m)$ is a mode-dependent initial amplitude, $r_{\text{LCFS}}$ is the distance between the coil and the LCFS, and $\delta r$ is a free parameter. Figure 14 shows the results of the fit of this function to the simulated perturbation amplitudes for a set of $n=0$ modes. For the fit, while $A_{0}$ is $m$-dependant and thus is fitted separately for each value of $m$, $\delta r$ is fitted simultaneously to all signals, ensuring it remains constant. As the multi-polar approximation is valid only when the measuring point is reasonably far from the plasma, the first coils have been omitted from the fit. The goodness of the fit for these few first coils improves with increasing mode number, but for $m=1$ notable deviations can be observed. The obtained value of $\delta r$ is consistent with a multipole located approximately over the central field conductor of the device, instead of being over the magnetic axis as expected. This is probably due to the shape of the plasma column, that differs notably from a cylinder, so some deviation from the ideal case is to be expected.
IV.1.2 Helical array
@@
The previous discussion on mode polarization, that was somewhat academic for the poloidal array, becomes essential for the helical array, as having tri-axial coils forces us to choose an orientation along which to project the temporal derivative of the magnetic field vector in order to conduct the analysis. The design orientations of the coils shown in 2 are a good starting point, as they project the signal into, roughly, its radial, toroidal, and poloidal components; yielding acceptable results.
@@
However, we can do better Haskey et al. (2015) by taking into account the magnetic configuration. We can take the direction of the magnetic field at the closest point on the LCFS for each coil, that for typical low-pressure, low-current TJ-II plasmas is very close to the (normalized) dual $\varphi$ basis vectorD’haeseleer (2012):
@@
and their binormal, that turns out to be the normalized tangent $\theta$ basis vectorD’haeseleer (2012):
$\hat{\mathbf{b}}\times\hat{\mathbf{n}}\simeq\hat{\mathbf{e}}_{\theta}$ (14)
@@
Figure 17: Lomb periodograms for three different modes ($m=2$, $m=4$, and $m=5$) calculated only with the $\hat{\mathbf{e}}_{\theta}$ projection of the signal, using only the coils in the helical array.
@@
$m_{\text{al}}=\dfrac{n_{\text{al}}}{N_{\text{fp}}}+l=\dfrac{n_{\text{al}}}{N_{\text{fp}}}+\dfrac{m_{\text{in}}N_{\text{fp}}-n_{\text{in}}}{N_{\text{fp}}}$ (15)
Figure 19: Same as figure18 for high mode numbers
@@ IV.2 Coupled modes simulations
Generally, Alfvén eigenmodes are destabilized in gaps in the Alfvén continuum that emerge from periodicities in the Alfvén velocity produced in turn by periodicities of the magnetic field structure. While, generally, perturbations in the plasma equilibrium are strongly damped if their frequencies lie on the continuum, such damping does not happen (or is much weaker) if their frequency falls in a gap. Such gaps can originate either from zeros in the radial variation of the continuum frequency ($\partial\omega/\partial r=0$) or from frequency crossings of counter-propagating waves Heidbrink (2008). For the latter, two (or more) modes with different mode numbers couple with a frequency that falls in the center of the gap. Depending on the mode numbers of the coupled modes, the AEs can be classified according to the symmetry that originates them (see the introduction to this section).
@@
For this scan, we introduce a quantity $\Delta$ that measures the separation between the radial maximum of the modes, so that, if $s_{0}=0.8$, one mode will be centered around $s_{1}=s_{0}-\Delta$ while the other will be centered around $s_{2}=s_{0}+\Delta$. Performing a scan over this quantity, and analising the results with the Lomb periodogram, we get figure 21. There, we see that the outermost mode is the only one that is identified, and that the periodogram intensity, that is an indication of the confidence in the identification, decreases when the modes overlap the most. Still, for $\Delta=0$, the identified mode is the $m=7,\ n=-11$ (although a considerably fainter maximum is also present for the $m=9,\ n=-15$ mode as well). The analysis shows that the diagnostic cannot resolve an HAE21 gap mode into its coupled modes, and that for modes with very close maxima, the component with lowest m (therefore higher signal as was shown in 14) will appear with higher intensity in the analysis. Whether or not this conclusion applies for TAEs (same $n$) or for modes with lower mode numbers is left for future analysis that can also benefit from experimental input.
V EfFect of plasma response on the determination of the poloidal mode number
All the results of the previous section have been obtained using a Biot-Savart integral in vacuum, without taking into account the plasma response.
The model described in A.2 can only account for the effect that the response of a homogeneous plasma of a given density immersed in a constant magnetic field would have on an arbitrary distribution of currents, and ignores the inconsistency of actually using a current distribution associated with a non-axisymmetric spatially extended magnetic field configuration. In order to obtain a meaningful estimate of this effect while minimizing the impact of approximating the response of the real non homogeneous plasma to that of a constant density one with magnetic field directed along the main direction of the real 3D field, we assume a highly localized mode current distribution centered toroidally at the toroidal plane where the poloidal array of Mirnov coils is located. Both the extended and toroidally localized potential distributions are shown in figure 23. This approximation is valid as long as we restrict ourselves to $n=0$ modes. Only in this case the synthetic signals generated by the localized current distribution are a good approximation to the ones generated by a toroidally extended realistic distribution of currents. This is illustrated in figure 24, where the corresponding synthetic signals are compared. Only slight differences appear and the $m=3$ structure is well preserved in the case of the localized perturbation.
Figure 23: On the left, an $n=0,m=3$ toroidally extended perturbation, which actually spans the whole torus although only a "straight" plasma column is shown. The same perturbation, radially localized around the detection plane of the poloidal array of Mirnov coils is shown on the right.
Figure 24: Synthetic signals generated by a toroidally extended (a) and a localized (b) $n=0,m=3$ perturbation.
To estimate the effect of the plasma response on the determination of the poloidal mode number, the localized distribution of currents, initially defined in the nodes of the boozer coordinates grid, is interpolated in a cartesian grid. For convenience, the grid and the current source are then rotated so that the $\mathbf{x}$ direction of the cartesian grid is perpendicular to the plane of the poloidal array, which lies now in the plane formed by vectors $\mathbf{y},\mathbf{z}$ of the grid. The 3D Fourier transform of the current source in this grid, $\delta\mathbf{J}_{ext}(\mathbf{r})$, is computed numerically using a multidimensional FFT algorithm and used in equation 40. Since we want to highlight the effect of the plasma response, the field is only evaluated at those sensors of the poloidal array whose spacelike distribution most closely matches the shape of the plasma (see white dots in figures 26 and 27).
We first perform a sanity check without plasma response to ensure that the magnetic field calculated at the coil locations using the Biot-Savart integral over the cartesian grid and the solution 42 based on Fourier decomposition methods yield similar results both in phase and wave amplitude. Figure 25 shows the normalized synthetic signals obtained with both methods and the wave amplitudes measured by each coil. Also, figure 26 shows the real part of the components of the perturbed magnetic field calculated at $t=0$ using the Fourier method, for $n=0,m=3$ and $n=0,m=9$ modes. The magnetic field perturbation that arises in response to the mode currents shows a variation of three orders of magnitude between the inner part of the plasma (milliTeslas) and the part where the coils are located (microTeslas) so we have chosen to represent the signed logarithm of the normalized values ($-\text{sign}(x)*\log(x)$). This choice of representation emphasizes the structure of the magnetic field perturbation outside the plasma, which is our region of interest.
As discussed in the appendix, solving the equation in Fourier space allows us to describe the plasma response through the dispersion tensor of each plane wave component. By taking a constant density and a constant static field $B_{0}$ along $\mathbf{x}$ in all the space domain we can take into account, approximately, the fast mobility of plasma electrons along the direction of the field and the subsequent shielding that they produce. Figure 27 shows the result of solving equation 31 considering now the contribution of the plasma response to the dispersion tensor given in 39. A homogeneous plasma with $n_{e}=10^{19}\ \text{m}^{-3}$, immersed in a $B_{0}=1$ T field, is taken to explore this effect.
Figure 26: Normalized components of the perturbed magnetic field in the central plane of the simulation grid $x=0$ for $m=3$ (left) and $m=9$ (right). White circles indicate the location of the poloidal coils. By construction, the coils are sentitive only to $\delta B_{z}$ and $\delta B_{y}$.
Evolving in time the field perturbation, for each of the cases represented in figures 26 and 27, allows us to calculate the synthetic signals and apply the Lomb periodogram analysis. Figure 28 shows the the results of these analysis. As expected from the result anticipated in figure 25, the mode structure without plasma response is well diagnosed by the set of coils for all the mode numbers considered. On the contrary, the plasma response clearly modifies the field structure and consequently the mode number detected by the coils differs from the input ones, the differences $\Delta m\equiv m_{out}-m_{in}$ reaching values up to $\Delta m=2$. This can be visually observed in figures 26 and 27 where it is shown that the structure of maxima and minima of $\delta\mathbf{B}$ is modified by the response of the plasma, in particular for those components of the magnetic field that have an impact on the coils measurement ($\delta B_{z}$ and $\delta B_{y}$). Note that the component of the perturbed field along the direction of the static homogeneous field ($\delta B_{x}$), which is not detected by the coils, is precisely the component less modified by the plasma response since mode currents flow mainly in the $\mathbf{x}$ direction and thus barely contribute to changes in $\delta B_{x}$.
Despite the approximations made in the plasma response model (i.e. modes with n=0 and toroidally localized current source), we can assume that it captures the fundamental aspects of the problem. Indeed, the detected mode numbers show differences with respect to the initial ones. However, this can only serve as an indication that the plasma has a clear effect on mode detection, but we cannot provide a clear quantification of this since the price of assuming an homogeneous plasma is that the electric field associated with the mode produces currents in the whole volume and not only in the part where the plasma is actually present. This can only be addressed with a full-wave code that solves this problem in real space.

### v3

@@ IV.1.2 Wave polarization
Figure 11: Ratio of squared azimuthal $|\delta B_{\zeta}|$ to radial-axial $|\delta B_{\text{RZ}}|$ amplitudes detected by the poloidal coils for $n=0$ modes with different poloidal mode numbers.
@@
Figure 12: 2D projection $\delta B_{\text{RZ}}(t)$ of the general $\delta{\mathbf{B}}(t)$ polarization ellipses for an $m=2$, $n=0$ mode. The perpendicular vectors to each coil winding plane ($\mathbf{P}$, blue arrows) and the vectors indicating the major (black arrows) and minor (red arrows) axis of the polarization ellipses at each location are shown.
@@
As a first approach, and considering the geometry of the poloidal array, we restrict ourselves to the projection of the general 3D polarization ellipse on the RZ plane, trying to find out if there is an optimal orientation for the coils, always keeping $\mathbf{P}$ in the RZ plane. In figure 11 we show the ratio between the amplitude of the oscillating field in the azimuthal direction ($|\delta B_{\zeta}|$) to that in the RZ plane ($|\delta B_{\text{RZ}}|$). A very low value of this ratio indicates that the oscillating field lies mostly in the RZ plane, so we can expect a better performance of the diagnostic in such cases. This ratio depends on both the mode numbers and the radial position of the modes, and there are instances where it is not small. Figure 12 shows the (normalized) projection of the $\delta B_{\text{RZ}}(t)$ polarization ellipse over the RZ plane with their major and minor axes and the vectors $\mathbf{P}$, that determine the coil orientations, for an $m=2$, $n=0$ mode. Note the changes in angle between $\mathbf{P}$ and the major axis of the ellipse, that become more pronounced on the rightmost coils.
@@
It might seem, after inspection of the result shown in figure 12, that mode identification accuracy could be improved by performing the periodogram analysis on the components of the field perturbation in the local polarization basis defined by the major and minor axes of the projected ellipses. That is, rotating $\mathbf{P}$ so that it aligns with the major/minor axis of the ellipse at each coil. However, this method has been tested and it has proved to not be effective, and projecting over the design orientation given by $\mathbf{P}$ (blue arrows in figure 12) yields much better results. Furthermore, as the polarization of the measured signal is mode-dependent, one cannot design an uni-axial coil array that performs well for any mode, and the experimental use of the aforementioned procedure is not feasible without having tri-axial sensors.
@@
Going back to figure 12, we see that the coils where the angle between $\mathbf{P}$ and the major axis of the ellipse changes more rapidly, when moving from coil to coil, are the ones that are closest to the plasma. Figure 13 shows the results of the mode analysis with the rightmost five coils removed on each side, hoping for a decrease in the distortion of the mode structure and an improvement of the detection accuracy. We observe that the detection accuracy does not improve, and off-by-one errors appear more frequently. This could be caused by a lack of angular coverage of the plasma column, but that would mainly affect the detection of low-$m$ modes ($m\leq 3$), which does not seem to be the case. Alternatively, the issue may lie in the coils furthest from the plasma, as the part of the plasma that contributes to the Biot-Savart integral with significant intensity increases with coil distance (as the contribution of a volume element is proportional to $1/r^{2}$). On the other hand, the detection limits (the radial position and mode width where the mode identification is not possible) do not change significantly.
@@
Finding a suitable polarization basis is essential for the helical array, as having tri-axial coils forces us to choose an orientation along which to project the temporal derivative of the magnetic field vector in order to conduct the analysis. The design orientations of the coils shown in 2 are a good starting point, as they project the signal into, roughly, its radial, toroidal, and poloidal components; yielding acceptable results.
@@
However, we can do better Haskey et al. (2015) by taking into account the magnetic configuration. We can take the direction of the magnetic field at the closest point on the LCFS for each coil, that for typical low-pressure, low-current TJ-II plasmas is very close to the (normalized) dual $\varphi$ basis vectorD’haeseleer et al. (1991):
@@
and their binormal, that turns out to be the normalized tangent $\theta$ basis vectorD’haeseleer et al. (1991):
$\hat{\mathbf{b}}\times\hat{\mathbf{n}}\simeq\hat{\mathbf{e}}_{\theta}\,.$ (12)
@@
IV.1.3 Helical array
Figure 16: Lomb periodograms for three different modes ($m=2$, $m=4$, and $m=6$) calculated only with the $\hat{\mathbf{e}}_{\theta}$ projection of the signal, using only the coils in the helical array.
@@ IV.1.3 Helical array
$m_{\text{al}}=\dfrac{n_{\text{al}}}{N_{\text{fp}}}+l=\dfrac{n_{\text{al}}}{N_{\text{fp}}}+\dfrac{m_{\text{in}}N_{\text{fp}}-n_{\text{in}}}{N_{\text{fp}}}\,.$ (13)
Figure 18: Same as figure 17 for high mode numbers Differences between input ($m_{\text{in}},n_{\text{in}}$) and output ($m_{\text{out}},n_{\text{out}}$) poloidal (a) and toroidal (b) mode numbers.
@@ IV.2 Coupled modes simulations
Generally, Alfvén eigenmodes are destabilized in gaps in the Alfvén continuum that emerge from periodicities in the Alfvén velocity produced in turn by periodicities of the magnetic field structure. While, generally, perturbations in the plasma equilibrium are strongly damped if their frequencies lie on the continuum, such damping does not happen (or is much weaker) if their frequency falls in a gap. Such gaps can originate either from zeros in the radial variation of the continuum frequency ($\partial\omega/\partial r=0$) or from frequency crossings of counter-propagating waves Heidbrink (2008). Similarly to what happens in tokamaks, where modes of the same $n$ and different $m$ can evolve jointly at the same frequency within the TAE gap, the existence of helicity gaps (HAE) in stellarators allows for complex radially extended and weakly damped structures characterized by modes with different values of $n$ and $m$. For instance, in the case of an HAEμν gap created by the appropriate components in the Fourier expansions of the magnetic field and the metric coefficients, modes with $n_{1}$ and $m_{1}$ can couple with modes with $n_{2}=n_{1}+\nu N_{\text{fp}}$ and $m_{2}=m_{1}+\mu$.
@@
For this scan, we introduce a quantity $\Delta$ that measures the separation between the radial maximum of the modes, so that, if $s_{0}=0.8$, one mode will be centered around $s_{1}=s_{0}-\Delta$ while the other will be centered around $s_{2}=s_{0}+\Delta$. Performing a scan over this quantity, and analising the results with the Lomb periodogram, we get figure 21. There, we see that the outermost mode is the only one that is identified, and that the periodogram intensity, that is an indication of the confidence in the identification, decreases when the modes overlap the most. Still, for $\Delta=0$, the identified mode is the $m=7,\ n=-11$ (although a considerably fainter maximum is also present for the $m=9,\ n=-15$ mode as well). The analysis shows that the diagnostic cannot resolve an HAE21 gap mode into its coupled modes, and that for modes with very close maxima, the component with lowest m (therefore higher signal as was shown in fig. 34) will appear with higher intensity in the analysis. Whether or not this conclusion applies for TAEs (same $n$) or for modes with lower mode numbers is left for future analysis that can also benefit from experimental input.
V Experimental results
To illustrate the usefulness of the synthetic diagnostic when analyzing real experimental data, we take an example from the TJ-II database. The chosen shot belongs to a series of experiments aiming at characterizing the influence of Electron Cyclotron Current Drive (ECCD) on the Alfvén spectrum of NBI-heated plasmas Cappa et al. (2021); Pons-Villalonga et al. (2023). The plasma was started with second harmonic ECRH, using two 53.2 GHz gyrotrons, with one of the ECRH beams configured to provide co-ECCD at lower power, helping also with density control. At $t=1125$ ms, co-NBI heating is introduced, and one of the gyrotrons is switched off shortly after.
Figure 22: Spectrogram of magnetic fluctuations (a), heating scheme (b), line integrated density (c), and plasma current (d) A red rectangle (top) illustrates the time-frequency range used to make figure 23.
The spectrogram of the magnetic fluctuations detected with one of the Mirnov coils is shown in figure 22-a). Below, the heating scheme (fig. 22-b), the line-integrated density (fig. 22-c) and the plasma current (fig. 22-d) of the discharge are pictured. The spectrogram shows that, a few milliseconds after turning on the neutral beam, a mode with $f_{1}\sim 320$ kHz, probably TAE or HAE appears. This mode presents a slight chirping character, mainly during the final phase of the discharge, where another chirping mode with frequency $f_{2}\sim 290$ kHz appears, at the same time than a low-frequency non-Alfvénic mode.
The changes in the MHD spectrum can be explained by changes in the rotational transform due to the stabilization of the NBI-driven part of the plasma current. This has been discussed elsewhere Cappa et al. (2023) and the physics interpretation as well as a precise mode identification analysis lies outside of the scope of this paper.
Figure 23: DMUSIC spectrogram of selected range.
This shot has been selected because the mode frequency changes minimally during its lifetime, thanks to a very good density control. This is helpful when conducting the analysis, as the Lomb periodogram is only valid for a constant mode frequency. In modes with relatively slow frequency changes, however, this limitation can be overcome by a piecewise analysis. In any case, a constant frequency is clearer for the illustrative purposes of this section. Figure 23 shows a spectrogram of a suitable time interval of the discharge, calculated with the DMUSIC algorithm Kleiber et al. (2021), where the narrow, constant-frequency nature of the mode can be appreciated.
Figure 24: Lomb periodograms of the $f_{1}\sim 320$ kHz mode, taken at three distinct times (rows). The left column shows the periodogram calculated using only the poloidal array signals. The middle column presents the same calculation but with the helical array. The rightmost column shows the product of the previous two.
The results of the analysis in three distinct intervals of the discharge of the $f_{1}\sim 320$ kHz mode are shown in figure 24. The results are very reproducible, although they lack the clarity of the results obtained using synthetic signals and an unambiguous identification of the dominant mode number is difficult, as the maximum in the product of periodograms (shown in the rightmost column) changes between $m=3,\ n=7$ in the first two rows and $m=1,\ n=-1$ in the last. This uncertainty can be attributed to the conditions of a real experiment where some of the sensors of the poloidal array, damaged by ECRH stray radiation, were not in operation.
Figure 25: Zoomed in spectrogram and identified mode numbers for the $f_{1}\sim 315$ kHz mode.
A time-resolved mode number measurement can be carried out by repeating this process for a series of suitable time intervals. In figure 25 the measured mode numbers obtained in this way are plotted over the zoomed-in spectrogram. We observe that for $t<1165$ the mode numbers are most commonly $n/m=7/3$, while for $t>1165$ the pair of mode numbers $n/m=-1/1$ are measured. To understand this result, or at least, to know to what extent we can take it as good, we used a combination of HIBP measurements and results provided by the synthetic diagnostic. Note that while the density remains constant along the entire length of the shot, the current does not. The combined effect of the Neutral Beam Current Drive (NBCD) and ECCD, that both induce positive current, results in an increase of the rotational transform which is here the main driver of changes in the spectrum. The radial profiles of the perturbed electrostatic potential, measured with one of the heavy ion beam probes are shown in figure 26 for $t<1162$ (a, top) and $t>1162$ (b, bottom). They exhibit a strong ballooning character typical of gap mode and, most importantly, the maximum of the profile appears to move inward for the second part of the shot. This is seen more clearly on the low field side ($\rho>0$) since the mode amplitude is larger there. Taking into account the synthetic diagnostic results regarding mode number errors when mode radial location shifts inward (17), the $/1$ 1 mode number measurement is very likely not accurate, corresponding with the unidentifiable core-localised modes that were discussed in the previous section. This underscores the synthetic diagnostic’s ability to effectively discern trustworthy mode number measurements from those that are not. However, this is not the only possible explanation. Alternatetively, changes in rotational transform could also explain why the destabilized modes are different, in addition to having different radial positions. To rule out this possibility, a more comprehensive set of simulations, out of the scope of the paper, is needed to precisely explain the behavior of this eigenmode.
Figure 26: Electrostatic perturbed potential profiles measured with HIBP, for $t<1162$ ms (a) and $t>1162$ ms (b). Vertical dashed lines mark the mean position of the mode during each time interval. $\rho<0$ corresponds to the high field side of the plasma, and $\rho>0$ to the low field side.
The above example illustrates how the synthetic diagnostic can help in the interpretation of the results but not simulate the measurement itself since it shows that, in the case of gap modes, the various coupled modes cannot be resolved to begin with. Therefore, in the above case, there is insufficient experimental information available to model the measurement of the coils.
The simulations and analysis conducted up until now have always been performed considering the best-case scenario: an exact positioning of the coils, no instrumental effects on the signal, no electromagnetic noise from the bulk plasma, no eddy currents in the nearby vessel nor in the metallic tubes housing the coils, etc. When considering experimental data, however, these sources of error must be taken into account. As a first approach to understand the impact that all these sources of error have on the experimental results, we can explore the effect that numerical noise added to the synthetic signals has on the mode analysis results. We take, for this excercise, a pair of mode numbers compatible with the ones detected experimentally. In figure 27, three periodograms made with synthetic data over the experimentally available coils of the poloidal array are shown. The first one, calculated with no noise, serves as a benchmark. Comparing with i.e. figure 8, we can already see a slight broadening of the band, that is caused by the decrease in the number of available signals.
Figure 27: Periodograms of a synthetic $m=1,\ n=1$ mode, using the poloidal array only. Left, with the noiseless signal. Center, with added gaussian noise. Right, with phase noise.
For the second one, gaussian noise has been added to the signals, according to
$y_{ij}^{\text{n}}=y_{ij}+\epsilon_{ij}\,$ (14)
where $\epsilon_{ij}$ is a random sample from a normal distribution with standard deviation scaled to, in this case, a 15% of the signal amplitude. The effects of this random noise are barely noticeable. We can think of the periodogram as a least-squares fit, and so intuitively this kind of noise should "cancel out" for a long enough time interval.
The third periodogram is the most interesting. There, a small random phase offset $\delta_{j}$ is added to the signal of each coil, via:
[…]
Here, $y_{ij}^{\text{c}}$ is the complex output of the synthetic diagnostic. This random phase offset models both any uncertainties in the spatial position of the coils, that would change their associated Boozer angles; the effects of eddy coils in the vaccuum vessel and nearby plasma-facing components, and the data acquisition system could have on the digitized signal. These have been observed and corrected in other devices Horváth et al. (2015) but a similar work has not been done yet in TJ-II. This phase noise significantly degrades the results of the periodogram, but if we assume it is either small (not necessarily the case, see ref. 45 again) or similar for all coils (more likely, as the surrounding environment and data acquisition pipeline are comparable), the mode number can still be identified.